A stone is dropped from a certain height which can reach the ground in 5 second. If the stone is stopped after 3 seconds of its fall and then allowed to fall again, then the time taken by the stone to reach the ground for the remaining distance is


        (a)    2 sec


        (b)    3 sec


        (c)    4 sec


        (d)    none of these


Grade:upto college level

1 Answers

Radhika Batra
10 years ago

(c)

    h = 1/2 g(5)2, h1 = 1/2 g(3)2

    h – h1 = 1/2 gt2

    1/2 g(5)2 – 1/2 g(3)2 = 1/2 gt2 ? t = 4 sec.